OpenStudy (anonymous):

find the GCF of 18x3 and 30x5

OpenStudy (anonymous):

can anyone help me please?

OpenStudy (anonymous):

if you decompose both numbers in factors, you'll get 18 = 2*3*3 and 30 = 3*2*5 so you now have 2*3*3*3 and 2*3*5*5, which means your GCF is 2*3, which is 6